OpenStudy (anonymous):

The equation of line LM is y = 5x + 4. Write an equation of a line perpendicular to line LM in slope-intercept form that contains point (-3, 2).

OpenStudy (anonymous):

Select one: a. y = -5x + 13 b. y=-1/5x + 13 c. y= 1/5 x - 7/5 d. y= 5x - 17

OpenStudy (anonymous):

@kc_kennylau

OpenStudy (kc_kennylau):

What's the slope of line LM? :)

OpenStudy (anonymous):

its positive

OpenStudy (kc_kennylau):

What exactly is the slope of line LM? :)

OpenStudy (nincompoop):

positive ... dohhhhh!! laughing out loud

OpenStudy (nincompoop):

y = mx + b is the intercept-form of an equation of a line, where m is your slope and b is the y-intercept

OpenStudy (anonymous):

ok soo.. what do I do now how am I supposed to find the answer

OpenStudy (nincompoop):

what you do now is figure out your damned slope

OpenStudy (anonymous):

ok how do I do that

OpenStudy (nincompoop):

have you been paying attention to our responses?

OpenStudy (nincompoop):

I will type it again y = mx + b is the intercept-form of an equation of a line, where m is your slope and b is the y-intercept

OpenStudy (kc_kennylau):

slope-intercept form*

OpenStudy (nincompoop):

an example of that crap slope-intercept form is y = 3x + 4 3 would be the slope and 4 would be the y-intercept form

OpenStudy (anonymous):

i know now ok let me see....

OpenStudy (nincompoop):

now, can you figure out your damned slope in the equation you were given in your problem?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

ok what would be m

OpenStudy (nincompoop):

so what on earth is the value of m in your problem?

OpenStudy (anonymous):

that's what im asking you

OpenStudy (nincompoop):

so you haven't been paying attention I am not here to answer your question, I am here to ask you the proper questions so that you can learn your lesson

OpenStudy (anonymous):

i put this question up so someone could help me figure out the answer not give me a math lesson

OpenStudy (nincompoop):

the purpose of this website is to guide people learn their materials, not be an accomplice to cheating

OpenStudy (anonymous):

was it implied that i wanted the answer

OpenStudy (nincompoop):

yah

OpenStudy (nincompoop):

y = mx + b is the slope-intercept form of an equation of a line, where m is your slope and b is the y-intercept an example of that crap slope-intercept form is y = 3x + 4 3 would be the slope and 4 would be the y-intercept form

OpenStudy (uh60blackhawk):

we are here to work with you not to just give out the answers.

OpenStudy (nincompoop):

giving you the math lesson is the guide to the answer you seek :) sorry we don't spoon feed answers like you're a helpless being.

OpenStudy (anonymous):

im not helpless im passing this class i just needed help on one question you don't have to help me

OpenStudy (nincompoop):

based on the definition and example, what the damned slope in your problem?

OpenStudy (nincompoop):

I think you're just going to accept defeat and not learn this overall